# Generative Large Language Models in Automated Fact-Checking: A Survey

### v1

@@ Abstract
The dissemination of false information across online platforms poses a serious societal challenge, necessitating robust measures for information verification. While manual fact-checking efforts are still instrumental, the growing volume of false information requires automated methods. Large language models (LLMs) offer promising opportunities to assist fact-checkers, leveraging LLM’s extensive knowledge and robust reasoning capabilities. In this survey paper, we investigate the utilization of generative LLMs in the realm of fact-checking, illustrating various approaches that have been employed and techniques for prompting or fine-tuning LLMs. By providing an overview of existing approaches, this survey aims to improve the understanding of utilizing LLMs in fact-checking and to facilitate further progress in LLMs’ involvement in this process.
@@ 1 Introduction
The contemporary digital era has brought various challenges, including the spread of false information. The prevalence of social media has made this problem more challenging and a serious societal concern. An effective strategy for combating false information involves fact-checking Vlachos and Riedel (2014), mostly done manually by fact-checkers. However, the number of fact-checkers is limited, and the spread of false information is increasing Aïmeur et al. (2023). Hence, there is an effort to develop an automated fact-checking process by employing digital tools and LLMs that would assist fact-checkers Nakov et al. (2021a).
LLMs utilize large-scale datasets, paired with billions of parameters, to model language nuances and patterns of natural language. Furthermore, generative LLMs represent a specialized subset of LLMs tailored for text generation. Their application in manual fact-checking has demonstrated promising results in enhancing the efficiency and accuracy of fact-checking. With this study, we provide the first comprehensive summarization of approaches and techniques that incorporate generative LLMs in the fact-checking process.
Existing surveys have explored the needs of fact-checkers and tasks that can be automated Nakov et al. (2021a), task definitions, traditional methods, or BERT-like architectures Thorne and Vlachos (2018); Zeng et al. (2021). Moreover, several authors have examined existing datasets for fact-checking Guo et al. (2022). Another survey aimed to combat false information using LLMs and explore opportunities and challenges that LLMs bring Chen and Shu (2023b). However, this study did not explore methods used in detail, which presents an opportunity to address the role of generative LLMs in fact-checking more closely.
Our main contribution is to provide an overview of the utilization of generative LLMs in automated fact-checking. We present an overview of 70 papers with relevant approaches and novel prompting techniques for researchers who want to delve into LLM-aided information verification. We identify four main tasks and various proposed strategies to tackle them. Furthermore, we discuss future challenges and possible directions for utilizing LLMs in the information verification process.
@@ 2.1 Selection
To identify a set of relevant articles, we employed a manual search of papers from NLP conferences, including ACL, EACL, NAACL and EMNLP. Along with these known conferences, we utilized a keyword search using Google Scholar and ArXiv.
In addition, we extended the created set to include papers referenced in related works and those that cited any article from our retrieved list of relevant papers. Additionally, papers submitted to CheckThat! Shared tasks were also involved in our survey. Overall, we collected 70 papers that employ generative LLMs for fact-checking.
@@ 2.2 Scope of the Survey
The survey is intended for researchers with expertise in the domain and a fundamental understanding of methods used across the NLP field. Our objective is to present an overview of techniques employed for fact-checking using generative LLMs. While fact-checking has evolved to include various modalities, our study delves only into the domain of textual and tabular information. Unlike other surveys that describe fact-checking datasets in detail, our focus lies on approaches, techniques, and associated challenges leveraging generative LLMs.
3 Tasks and Approaches Employed in Fact-Checking
In this section, we delve into the particular tasks of automated fact-checking, describing its definitions and the methods employed when engaging generative LLMs in the fact-checking process.
3.1 Check-Worthy Claim Detection
The task of check-worthy claim detection concerns identifying claims that contain potentially misleading information and that should thus be checked and verified. Selecting those claims is challenging, and there is no clear definition of a check-worthy claim. However, claims that contain information that may be false or misunderstood and are potentially harmful to society are usually considered to be check-worthy claims.
Generally, this task is considered a binary classification, the most common approach when employing LLMs. However, a finer-grained classification can be used to determine the harmfulness of a claim Wang et al. (2023b). Additionally, several authors have investigated the generative capabilities of LLMs to generate claims, where LLMs generate a set of verifiable claims from a given text, or questions, which aims to produce questions for further verification Chen et al. (2022). The advantage of using generative abilities, specifically the above techniques is that LLMs are able to extract key information and decompose complex statements into easily verifiable parts or questions.
@@ 3.2 Previously Fact-Checked Claims Detection
In previously fact-checked claims detection, LLMs are usually used to rank the already retrieved fact-checks, e.g., retrieved using the similarity of text embeddings. One possible application is to classify pairwise relationships between claims and retrieved fact-checks as textual entailment. In addition, LLMs can be employed to filter out irrelevant fact-checks that have been retrieved using another ranking approach.
3.3 Evidence Retrieval
The objective of evidence retrieval is to gather crucial information from trusted sources to assess the claim’s veracity, which can have multiple forms, e.g., text or tables. The purpose of evidence is to provide the knowledge needed for fact-checking to assess the veracity of the claims that are being fact-checked.
When incorporating LLMs into the fact-checking, the evidence fills knowledge gaps in LLMs, keeping them updated with up-to-date information they lack. In our paper, we refer to this setup as open-book setting, which allows LLMs to access external knowledge, such as Wikipedia or open web Schlichtkrull et al. (2023). However, selecting the knowledge source is crucial, and Wikipedia and the open web have several weaknesses. Thus, scientific reports Bhatia et al. (2021) are a more credible choice. Conversely, some approaches rely only on internal LLM knowledge and are termed closed-book settings.
The most employed approach for evidence retrieval is ranking. Alternative strategies explored the generative abilities of LLMs for query generation, in which LLM generates a search query for web search, and rationale selection, in which the LLM is tasked to select relevant sentences out of the previously retrieved documents.
3.4 Fact Verification & Fake News Detection
Fact verification focuses on determining the truthfulness of a claim or text based on the previously retrieved evidence. On the other hand, the fake news detection tries to classify a given input without the evidence based solely on the provided input.
The most prominent approach employing generative LLMs involves classifying a claim into three categories (supports, refutes, or not enough information) based on the given evidence. However, it can also be considered a binary or more fine-grained classification problem, e.g., six Buchholz (2023) or seven categories Hoes et al. (2023a). In addition to classification, explanation generation is a commonly used technique when incorporating generative LLMs for fact verification, where the goal is to explain the LLM’s decision and clarify what is misleading on a given claim.
@@ 4 Taxonomy of Methods
Figure 1: A taxonomy of methods employed when incorporating generative LLMs in the fact-checking process.
Generative LLMs have been employed in fact-checking through several approaches, which we divided into three categories: Classification & Regression, Generation, and Synthetic Data Generation. These categories differ in the granularity of the LLM’s output. In Classification & Regression, LLMs categorize inputs into discrete labels or continuous values. In Generation, the output is continuous text based on the input data. For Synthetic Data Generation, LLMs generate parts or entire datasets for training, validating and testing fact-checking methods. The proposed taxonomy is illustrated in Figure 1 and a list of relevant papers and detailed information is in Tables 2, 3, 4 and 4.
Classification & Regression and Generation are further categorized into prompting, fine-tuning and augmentation with external knowledge. Prompting leverages instructions to guide responses without LLM’s adaption, fine-tuning adapts LLMs to specific tasks via training on target tasks and knowledge augmentation enhances the responses of LLMs by incorporating up-to-date information.
4.1 Classification & Regression
In classification, LLMs predict categorical labels, while in regression LLMs are instructed to provide continuous values as predicted scores on various scales (0-1, 0-100, etc.). However, regression is often reframed into a soft classification problem due to the absence of datasets with score targets. In soft classification, the authors leverage the produced scores to determine the final class using thresholds (commonly 50%). The benefit of soft classification is that the threshold for the classes can be optimized based on the identified properties of LLMs.
Since the objective of most fact-checking tasks is to classify a claim into a certain category, classification is the most commonly used method in fact-checking, especially in check-worthy claim detection (e.g., yes or no) and fact verification (e.g., false, half true, true). Nevertheless, its potential use is also in other tasks, such as detecting previously fact-checked claims as an entailment classification between a given statement and previously fact-checked claims. In contrast, regression is widely employed in previously fact-checked claims detection and evidence retrieval, where the identified claims or evidence are ranked using a predicted score. Furthermore, regression is leveraged to a small extent in fact verification, intending to predict the truthfulness score for a given claim.
4.1.1 Prompting
In the era of generative LLMs, prompting has gained prominence, which is a method of instructing the LLM to condition the response to achieve the desired result. This method is employed in many NLP tasks, and several authors investigated the capabilities of prompting in the fact-checking domain. In classification, prompts with a list of possible categories are commonly used, while LLMs are instructed to choose the most probable label. These prompting techniques are particularly beneficial when datasets are limited across various languages or domains, making traditional fine-tuning methods challenging to use. This is especially a concern for fact-checking datasets which are limited mostly by their size.
Zero-Shot Prompting.
In zero-shot prompting, LLMs are directly prompted with the instruction without providing examples. This is particularly beneficial when there is a lack of extensive training data and quality samples to serve as in-context examples. However, zero-shot prompting imposes several limitations. One of them is that instructions need to be clear and describe the task context to perform it correctly. It is crucial to introduce all the terminology within the prompt to mitigate ambiguities, e.g., attributes of check-worthy claims. Due to these challenges, prompt engineering is an essential part of experiments that need to be conducted, especially for fact-checking.
Another limitation is that the instruction wording and style can significantly affect LLM performance. To address this issue, various prompt styles have been proposed. In fact-checking, especially role specification Li and Zhai (2023) and JSON format Ma et al. (2023) have shown promise in enhancing performance. Role specification involves instructing LLM to act based on a defined role, such as fact-checker, while in the JSON format, the instruction adheres to JSON structure (e.g., Check the ${claim}\nAnswer:$). Role specification aids LLMs in better understanding the task, while the JSON facilitates easier post-processing due to its structured response format.
Since fact-checking poses a challenging task for zero-shot prompting and provided answers by LLMs may not always be based on facts, Zeng and Gao (2023) utilized multiple prompt variants to assess the veracity label of each claim. These prompt variants differ in only one word and represent various relationships between the claim and whether it is true, unclear or false, to avoid ambiguities in the output of the LLM. This technique verifies the consistency of LLMs’ responses to a given claim.
@@ 4.1.1 Prompting
Zero-shot prompting has one major limitation: the model is not always able to capture the essence of the task based only on the given instruction. Therefore, few-shot prompting is often used to enhance LLM performance by providing demonstrations necessary for complex tasks like fact-checking. LLMs efficiently learn the nuances of the task during inference utilizing provided labelled examples. Applying few-shot prompting is similar to zero-shot, whereas the authors also investigated the number of samples provided to the LLM Cao et al. (2023); Zeng and Gao (2023, 2024).
@@
The Chain-of-Thought (CoT) technique harnesses the power of LLMs to provide logical reasoning steps. This reasoning is performed before the actual class is provided, with the LLM exploiting its reasoning capabilities by incorporating intermediate reasoning steps for the final assessment. The CoT can be used in zero-shot Hu et al. (2023); Chen and Shu (2023a) and few-shot Liu et al. (2024a) settings, where authors include examples of the steps that reasoning should involve, e.g., follow-up questions Sawiński et al. (2023).
4.1.2 Fine-Tuning
Fine-tuning represents the method for adapting LLMs to a specific task or domain. It is mainly employed when the LLM does not have the necessary capabilities to address a given task using prompting, especially in cases where the LLM has no expertise in the selected domain or the task is complex. However, due to a large number of parameters, fine-tuning is usually only a concern for smaller LLMs. For larger LLMs, fine-tuning is often not feasible due to the limited resources, and also, it is not necessary, since these LLMs have improved capabilities for solving a wide range of tasks, even without further fine-tuning. In fact-checking, fine-tuning is primarily employed for ranking tasks, e.g., previously fact-checked claims detection and evidence retrieval. However, it is also used for other tasks beyond these applications.
LLM fine-tuning can be done in multiple settings and one of them is LLM fine-tuning on a single task. The advantage of such training is the specialization of the LLM to only a specific task and hence the improved ability of the LLM to address it. Conversely, the drawback of this approach lies in the fact that the LLM loses the ability to solve other tasks, even if they are from the same domain (e.g., fact-checking), known as catastrophic forgetting McCloskey and Cohen (1989). In this context, addressing multiple tasks necessitates fine-tuning a specific LLM for each task. One technique for LLM fine-tuning on a single task involves leveraging the logits of the particular token produced by the LLM, such as true token, which can be used for document ranking Pradeep et al. (2021). We refer to document ranking based on the probability of the true token as classification ranking because the LLM has been trained in a binary classification setting. Building upon classification ranking, multiple authors utilized additional steps Jiang et al. (2021); Pradeep et al. (2021); Fernández-Pichel et al. (2022). The classification ranking approach can be further employed to rank evidence in evidence retrieval and to extract pertinent sentences from retrieved evidence Pradeep et al. (2020).

### v2

@@ Abstract
The dissemination of false information on online platforms presents a serious societal challenge. While manual fact-checking remains crucial, Large Language Models (LLMs) offer promising opportunities to support fact-checkers with their vast knowledge and advanced reasoning capabilities. This survey explores the application of generative LLMs in fact-checking, highlighting various approaches and techniques for prompting or fine-tuning these models. By providing an overview of existing methods and their limitations, the survey aims to enhance the understanding of how LLMs can be used in fact-checking and to facilitate further progress in their integration into the fact-checking process.
@@ 1 Introduction
The modern digital era has introduced numerous challenges, including the widespread dissemination of false information, a problem exacerbated by the rise of social media. Fact-checking is a key strategy for combating false information Vlachos and Riedel (2014), but it is still largely carried out manually by fact-checkers. As the number of fact-checkers remains insufficient to keep pace with the growing volume of misinformation Aïmeur et al. (2023), efforts are underway to develop automated fact-checking systems. These systems leverage artificial intelligence and LLMs to reduce the work of fact-checkers Nakov et al. (2021a).
The application of LLMs in automated fact-checking has demonstrated significant potential for improving both the efficiency and accuracy of fact-checking Wang et al. (2023a). LLMs are trained on large-scale datasets, and they leverage billions of parameters to capture nuances and patterns of natural language. Moreover, generative LLMs, specifically designed for text generation, open new possibilities for integrating LLMs into the verification process by generating meaningful, natural text that was not possible before and automating complex reasoning tasks.
Previous surveys on fact-checking in NLP domain have primarily focused on the needs of fact-checkers Nakov et al. (2021a); Hrckova et al. (2024), specific fact-checking tasks Guo et al. (2022b), datasets Guo et al. (2022a) or traditional and BERT-based methods Thorne and Vlachos (2018); Zeng et al. (2021). One recent survey addresses combating misinformation using LLMs, illustrating both the opportunities and challenges they present Chen and Shu (2023b). However, this study lacks a detailed exploration of the methods employed. This presents a gap for a more thorough examination of generative LLMs’ role in fact-checking.
Our main contribution is a comprehensive overview of approaches and limitations in using generative LLMs for automated fact-checking. We review 69 papers, highlighting relevant methodologies and innovative prompting techniques for researchers exploring LLM-aided information verification. We analyze fact-checking tasks, LLM methods, common techniques and languages covered in surveyed papers. Additionally, we outline future challenges and potential directions for advancing the use of generative LLMs in the verification process.
@@ 2.1 Selection
To compile a set of relevant articles, we conducted a manual search of papers presented at ACL, EACL, NAACL and EMNLP conferences. Additionally, we performed a keyword search using Google Scholar and ArXiv. We expanded this collection by incorporating papers cited in related works and those that cited any article from our initial list. Furthermore, we included papers from the CheckThat! Shared tasks. In total, we gathered 69 papers that employ generative LLMs for fact-checking. Additional details about our selection process are provided in Appendix A.
2.2 Observed Aspects
In our survey, we categorized the papers based on several key aspects. First, we investigated fact-checking tasks addressed in each study (Section 3) and the methods employed to incorporate LLMs for tackling these tasks (Section 4). These methods represent a high-level abstraction of the approaches used, categorized based on the output type generated by LLMs. Another aspect we examined is techniques (Section 5), which refer to the strategies, such as prompting and fine-tuning, used to adapt and apply LLMs for specific fact-checking tasks to ensure they generate desired outputs. Given that false information is a global issue, we also analyzed languages covered in each study (Section 7). Furthermore, we classified what type of knowledge base was used for evidence retrieval (see Table 5) and whether LLMs utilized evidence for the fact-verification task (see Table 5). Relevant studies for each task are summarized in Tables 3, 4, 5 and 5.
@@ 2.3 Scope of the Survey
This survey is intended for researchers with expertise in NLP and a fundamental understanding of methods across the field. We aim to present an overview of techniques employed in fact-checking using generative LLMs. While fact-checking has evolved to cover multiple modalities, our study is confined to textual and tabular data analysis. Unlike surveys that provide detailed descriptions of fact-checking datasets, our focus is on the methods, techniques, and challenges of leveraging generative LLMs. Our aim is to focus on the verification of the truthfulness of a given input using LLMs, rather than on evaluating LLM outputs for hallucinations Ji et al. (2023) or factuality Wang et al. (2023a, 2024).
3 Fact-Checking Tasks
Fact-checking is the process of verifying the truthfulness of a given claim Guo et al. (2022b). This process involves several distinct tasks, each essential for determining the credibility of the information. In this section, we sort identified tasks based on the number of related papers and highlight the most common methods used for each task.
3.1 Fact Verification & Fake News Detection
Fact verification and fake news detection are the most prominent tasks when incorporating LLMs into the fact-checking process, with 53 out of 69 surveyed papers aimed at these tasks. Fact verification involves assessing the veracity of a given claim or few sentences. In contrast, fake news detection aims to check the trustworthiness of longer texts, such as news articles Li and Zhou (2020).
[…]
3.2 Evidence Retrieval
Thirteen research papers have addressed the task of evidence retrieval using LLMs. Evidence retrieval is about gathering essential information from trusted sources to assess the veracity of claims. The evidence can take various forms, such as text or tables. This evidence serves as the knowledge necessary for accurate fact-checking. The selection of an appropriate knowledge source is crucial, as Wikipedia and the open web exhibit several weaknesses. In contrast, scientific reports Bhatia et al. (2021) offer a more credible alternative.
[…]
Figure 1: A taxonomy of methods for integrating generative LLMs in fact-checking, illustrating examples of model inputs and outputs. The methods employed with generative LLMs are classified based on the output type.
3.3 Claim Detection
Eleven of 69 papers addressed the claim detection task, indicating a moderate emphasis. Claim detection concerns identifying claims containing verifiable information necessitating verification. These claims, often termed check-worthy or verifiable claims, can include statements that may be false, misleading, or potentially harmful to society.
[…]
@@ 3.4 Previously Fact-Checked Claims Detection
The task of previously-checked claims detection aims to reduce redundant work for fact-checkers. Currently, it is the least explored area using LLMs, with only five identified papers. However, it is essential for fact-checkers to ascertain whether a claim has been previously fact-checked, as misleading claims propagate across various sources and languages Pikuliak et al. (2023b). Thus, the goal of previously fact-checked claims detection is to identify similar claims from a database that have already undergone fact-checking.
@@ 4 Taxonomy of Methods
In this section, we discuss the methods of using LLMs to perform various fact-checking tasks. We categorize methods based on their output type. This taxonomy is illustrated in Figure 1: (1) Structured output is used when LLMs process samples and generate answers in a predefined structure: a class for a classification task, a number for a regression task, ranking of inputs, etc. In this case, LLMs often replace older specialized approaches. (2) Unstructured output is used when LLMs process samples and generate texts such as explanations, summaries, etc. that are used in the fact-checking process. (3) Synthetic data generation is an approach when LLMs generate new samples for our datasets. These datasets can then be used to train other (often smaller and specialized) models.
4.1 Structured Output
Methods that produce structured output are the most common in fact-checking, appearing in 59 out of 69 papers. These methods are advantageous because they allow easier evaluation, offer higher execution efficiency and easier fine-tuning due to the availability of datasets with labeled targets. This makes them particularly beneficial for fact-checking tasks. In this section, we further categorize these methods into three output types: classification, regression and ranking.
Classification.
LLM-based classification utilizes LLMs to predict categorical labels from a predefined set, e.g., veracity labels. In this approach, LLMs are tasked to select the most probable class. Classification is commonly employed in fact-checking (54 out of 69), as many tasks involve categorizing claims, especially in claim detection (yes/no) or fact verification (e.g., false, half true, true). It is also effective in detecting previously fact-checked claims, as an entailment classification between a given statement and previously fact-checked claims Choi and Ferrara (2023, 2024).
Regression.
[…]
Ranking.
[…]
4.2 Unstructured Output
Methods with unstructured outputs typically involve generating continuous text, which poses greater challenges in assessment. However, these methods provide more detailed responses, often including justifications, which enhance fact-checking by offering insights into the LLM’s reasoning. Since generative LLMs produce open-ended outputs rather than selecting from a set of labels, they can address a broader range of fact-checking tasks.
We categorized these methods based on the output type generated by LLMs into three categories: (1) coherent, meaningful text consisting of multiple sentences; (2) single-sentence outputs; and (3) outputs consisting of only a few keywords or phrases.
Explanation Generation.
[…]
In addition to directly prompting LLMs to explain their decision, summarization can also be used as a form of explanation generation, in tasks like fact-verification Russo et al. (2023b). However, the limitation of summarization lies in its tendency to omit crucial details from the retrieved evidence, leading to summaries that may not fully address misleading claims, nor be as persuasive as traditional fact-checking articles.
Claim and Question Generation.
The goal of claim generation is usually to produce one sentence that captures all key information for various tasks, including claim detection, evidence retrieval or fact-verification. The claim generation can also be known as claim normalization Sundriyal et al. (2023), where the aim is to extract a single normalized claim from a social media post. However, challenges arise when multiple claims are present in the input, complicating the creation of a single normalized claim. Therefore, claim generation can also produce a list of claims.
A part of claim generation is span detection Gangi Reddy et al. (2022), which aims to identify the exact boundaries of the claim within the text. The advantage of span detection lies in its ability to select the precise wording of the claim.
Claim generation and question generation share a similar output format – a single sentence. However, in question generation, the output specifically consists of a question, mostly used to verify a given claim Chen et al. (2022).
Query Generation.
Another form of unstructured output is query generation Prieto-Chavana et al. (2023), where the output comprises several words or short phrases primarily used for retrieving evidence from external sources, such as search engines. This method allows LLMs to extract essential parts of a claim, enhancing retrieval efficiency and minimizing irrelevant results that may arise from using the full claim.
Next hop prediction Malon (2021) is a specific type of query generation aimed at identifying and generating the title with a sentence of the evidence to be retrieved. This is especially important when previously retrieved evidence introduces new entities that require further elaboration.
4.3 Synthetic Data Generation
In synthetic data generation, LLMs are employed to create entire datasets or their parts. Unlike the generation from Section 4.2, where LLMs are used to generate results, synthetic data generation creates data that are then used for training or fine-tuning. This method proves valuable when datasets are either unavailable or limited for specific tasks or languages. In fact-checking, it can facilitate the creation of datasets containing disinformation Chen and Shu (2023a); Vykopal et al. (2024) or fake-news Jiang et al. (2023); Huang and Sun (2023).
Synthetic data generation can augment datasets by filling gaps in underrepresented categories or increasing variability. For instance, LLM can generate debunked claims to extend datasets Singh et al. (2023), produce emotional misinformation tweets Russo et al. (2023a) or create counterparts for claim-evidence pairs with contrasting veracity and high word overlap Zhang et al. (2024a).
5 Techniques
In this section, we analyze how different studies applied techniques to guide LLMs in achieving accurate outputs. We observed several popular techniques to improve the performance of LLM, which we categorized into three groups: prompting, fine-tuning and augmentation with external knowledge. These techniques can overlap, with some studies using combinations, such as linking prompts with external resources. Papers were classified by the most prominent technique; for example, if prompting was combined with external knowledge, they were categorized under augmentation with external knowledge.
5.1 Prompting
Prompting is a set of techniques used to improve the performance of LLMs by designing the instructions. This approach is versatile and particularly effective in scenarios with limited data, such as fact-checking, where datasets are often scarce and labeled data is difficult to obtain.
However, prompting faces several limitations, especially with complex tasks like in fact-checking. For instance, detecting previously fact-checked claims involves ranking, where prompting alone may struggle to produce accurate results Choi and Ferrara (2024). Another limitation is the necessity for clear instructions and contextual descriptions to ensure accurate task execution. To reduce ambiguities, it is essential to define all relevant terminology within the prompt, e.g., the attributes of check-worthy claims. Therefore, prompt engineering is important, and LLMs can assist in prompt creation Cao et al. (2023).
Prompt structure and wording can influence LLM performance. Techniques like role specification Li and Zhai (2023) and JSON formatting Ma et al. (2023) have been shown to enhance performance. Role specification assigns the LLM a defined role (e.g., fact-checker), improving task comprehension, while JSON formatting structures instructions and outputs in a consistent manner (e.g., Check the ${claim}\nAnswer:$). Additionally, output sequencing affect performance Vergho et al. (2024); Pelrine et al. (2023). Presenting the explanation first encourages LLM reasoning, leading to more accurate assessments, whereas placing the final verdict first may cause the LLM to focus on justifying the outcome rather than producing a truthful explanation.
Fact-checking is challenging for prompting LLMs, as their responses can be inaccurate. Zeng and Gao (2023) addressed this by using multiple prompt variants, differing by a single word, to represent different relationships to the claim (e.g., true, unclear, or false). This method helps ensure consistency in the LLM’s output, as small changes can affect predictions. Moreover, LLMs can be prompted to rewrite claims in a neutral style, improving evaluation reliability Wu and Hooi (2023) and resilience to adversarial attacks.
Prieto-Chavana et al. (2023) used prefixes like Search query: to guide LLMs in generating search queries for evidence retrieval. Furthermore, by combining query generation with rationale selection, relevant documents are retrieved, and then the most pertinent information is selected.
@@ 5.1 Prompting
[…]
In-context examples are particularly useful in methods such as span detection Gangi Reddy et al. (2022) or claim generation Kamoi et al. (2023b); Li et al. (2023a). To reduce false positives, it is crucial that the examples capture a broad range of characteristics typical for the given task. Decomposing input text into sub-claims helps break down complex statements into smaller, verifiable components. However, during this process, LLM may unintentionally reformulate claims, potentially altering their meaning or veracity labels. By providing varied demonstrations, few-shot prompting can aid LLMs in comprehending the variability and defining characteristics of check-worthy claims.
Few-shot prompting extends beyond claim detection. It is also employed in evidence retrieval, such as generating search queries Li et al. (2023a) or selecting relevant rationales Wang et al. (2023b).
@@
[…]
5.2 Fine-Tuning
[…]
Single-Task Fine-Tuning
One approach to fine-tuning involves fine-tuning on a single task Agrestia et al. (2022); Petroni et al. (2021); Bhatia et al. (2021); Hyben et al. (2023); Althabiti et al. (2023); Russo et al. (2023b); Zeng and Zubiaga (2024), which enhances the model’s specialization and performance in that area. However, this method may lead to the phenomenon of catastrophic forgetting, where LLMs lose the ability to perform other tasks, even those within the same domain Luo et al. (2024). Consequently, addressing multiple tasks requires fine-tuning a specific LLM for each task. A strategy for single-task fine-tuning involves classification ranking Pradeep et al. (2021). This approach has been extended in previous works, e.g., by including named entities within the claim Jiang et al. (2021), reranking of previously retrieved documents Pradeep et al. (2021) or predicting the reliability of the retrieved passage based on a given claim Fernández-Pichel et al. (2022). Classification ranking can be employed to rank evidence in evidence retrieval and to extract pertinent sentences from retrieved evidence Pradeep et al. (2020).
Multi-Task Fine-Tuning.